# TPIA: Towards Target-specific Prompt Injection Attack against Code-oriented LLMs

### v5

@@ V-A Main Settings
TABLE II: The ASR across different Code LLMs. The experiment targets four code completion datasets and tests the success rates of various attack methods using three Code LLMs. As shown in Table I, some datasets do not include data for the C/C++ and Java languages, thus we have a total of 7 subdatasets. Best attack results are marked in bold.
@@
Malicious Objectives and Code Datasets. We hereby collect 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) [1] across 3 programming languages: Python, Java, and C/C++. The straightforward objectives include using MODE_ECB for encryption (dubbed ‘ST0’), using ssl3 protocol (dubbed ‘ST1’), and executing “rm -rf /” (dubbed ‘ST2’). ST0 and ST1 are studied in classical code backdoor attacks [46], while ST2 is a famous destructive instruction. CWE vulnerabilities are generally selected for experiments in code model research [23, 56], and CWE Top-10 KEV weaknesses are supported by American cybersecurity and infrastructure security agency (CISA). We construct cases using the method described in Section IV-B based on the official examples (i.e., CWE Top-10 KEV) or LLM-generated demo (i.e., ST0, ST1, ST2). As shown in Table I, some objectives apply only to C/C++, e.g., CWE416 “Use After Free”, thus we have a total of 31 cases. In addition, we collect corresponding code completion datasets, including Humaneval [14], Humaneval-x [61], CodeXGLUE [38], MBPP [10], and Eval-Plus [35]. The items in these datasets contain prompts and corresponding outputs across different code completion tasks, which can simulate unpredictable users’ instructions and context in real-world scenarios. Some datasets do not cover Java or C/C++, as shown in Table I. As a result, we have 7 sub-datasets (i.e., the subsection of a dataset containing a particular programming language) containing more than 14,000 data items totally.
@@
Model Selection. In our experiment, we select three models from two representative Code LLM series, including CodeGemma [49] and Code-GeeX [61], for evaluation. CodeGemma series are trained by Google LLC and outperform some old-version Code LLMs like CodeLlama [44] on code completion tasks. We select CodeGemma-2b and CodeGemma-7b to compare the performance differences between models of distinct scales. CodeGeeX series have corresponding applications [2] with more than 535k downloads (on VScode [6] Apps shop). We select CodeGeeX2-6b, a version with stable performance. We exclude earlier small-scale code-oriented models, such as CodeBERT [18] and CodeT5 [51], because large-scale code-oriented models perform much better than them in code completion tasks.
@@
Baseline Selection. We compare our TPIA with GCG-based and manually designed IPI attacks. In particular, the vanilla GCG [63] is a representative PI attack and does not meet the requirements of our attack scenarios. Therefore, we adapt GCG into an IPI attack that is applicable to Code LLMs, dubbed ‘C-GCG’ (code-oriented GCG). The manually designed IPI method (dubbed ‘IPI-M’) is designed based on the existing work [22]. IPI-M limits the number of injected tokens to make the perturbation concise since the existing IPI attack usually uses a large number of readable comments. Besides, we also provide the results of the original model (dubbed ‘No Attack’) for reference, since the original model itself may also have inherent insecurity factor.
@@
where $r_{1}$ and $r_{2}$ are set as 20% and 80% in our evaluation.
@@
Hyperparameters and Computational Facilities. We conduct experiments using the following parameter settings to evaluate our approach. During gradient calculation, we utilize five noisy code samples, and the random seed for initializing perturbations is set to 100. For forward reasoning enhancement in Section IV-D, the number of tokens reinforced is set to $h=2$, as defined in Equation 6. For the hyperparameters related to attack performance, the default parameter values are as follows: the Top-$k$ value is 400, the rate ($r$ in Equation 6) is 0.4, the number of keywords is 2, and the perturbation length is 10. Tokenization is performed using Gemma’s tokenizer.
@@
We run Algorithm 1 on an NVIDIA RTX A6000 Graphics Card with the above parameter settings, employing a float16 precision format without quantization. In general, the perturbation generation process requires approximately 0.5-1 hour for CodeGemma-2b and 2-3 hours for both CodeGemma-7b and CodeGeeX-6b. These computational times ensure the generation of high-quality perturbations tailored to the respective model sizes across all the experiment threat cases.
V-B Evaluation on Open-source Code LLMs
@@
We first perform the attack evaluation over open-source Code LLMs. This includes three aspects: (1) we assess the attack effectiveness over all evaluated Code LLMs across all threat cases, demonstrating that TPIA is widely effective and more threatening than other attack methods; (2) we evaluate TPIA’s flexibility against different targets using two special metrics; (3) we evaluate the trade-off between effectiveness and stealthiness $w.r.t.$ the number of injected tokens.
@@
Effectiveness. We hereby evaluate whether TPIA can successfully induce the Code LLMs to generate attacker-specified targets. We control the number of injected tokens for three attack methods in this experiment. For TPIA, we use its default settings: perturbation length of 10 and 2 keywords, while C-GCG uses a prefix length of 12. Since IPI-M requires a sufficient number of tokens to be effective, we just limit its perturbation to fewer than 30 tokens. Table II provides the average ASR for four methods. The effectiveness of TPIA significantly outperforms the other baseline methods: it achieves an average ASR of over 80% across all three Code LLMs, with the highest average ASR reaching 86.6%.
@@
Comparing the three Code LLMs, the CodeGemma-7b model has the highest security in normal conditions (8.4% for No Attack) and is the least likely to generate vulnerable code. However, it is the most susceptible model to TPIA, which may be attributed to its strong contextual learning ability. Comparing the three programming languages, Python is the most susceptible language to TPIA due to its concise syntax and the smallest number of tokens consumed; Besides, TPIA is relatively less effective on CodeXGLUE, which mostly due to its larger data length and more complex data content. In contrast, the simplest dataset with the shortest data items, MBPP, is the most vulnerable to attacks, including C-GCG and IPI-M. Therefore, we believe that an increase in the contextual complexity can reduce the ASR of TPIA. In other words, when the contextual code becomes complex, the strength of TPIA (e.g., the length of perturbations) needs to be increased accordingly to ensure the attack’s effectiveness. According to our subsequent experiments in Section V-C and Section V-D, increasing the length of TPIA or the Top-$k$ value can effectively enhance the attack performance of TPIA.
@@
TABLE III: The attack evaluation results of NBR and STF against open-source Code LLMs.
Method Metric C/C++ Python Java
  No Attack NBR 0.00 0.22 0.01
STF 0.00 0.11 0.03
C-GCG NBR 0.39 0.50 0.56
STF 0.26 0.47 0.39
IPI-M NBR 0.08 0.25 0.17
STF 0.09 0.19 0.07
TPIA (ours) NBR 0.77 0.92 0.78
STF 0.76 0.87 0.74
@@
Flexibility. The previous experiments show that TPIA has satisfactory effectiveness across all threat cases. However, it is also important to consider the flexibility requirement in practice. In general, it can be divided into two situations. One is when attackers want to achieve diverse malicious objectives, which requires that cases with low ASR be minimized. The other is when attackers want to cause severe damage, which requires that cases with high ASR be maximized. We use NBR to evaluate the attack flexibility for the first situation, and STF for the second situation.
@@
As shown in Table III, TPIA has the highest NBR on the Python programming language, which means that implementing viable threat cases in Python is the easiest, and attacks targeting Python files are more flexible. Additionally, TPIA scores the highest in STF for Python, indicating that the attack effects in Python are more significant and pose a more serious real-world threat. STF and NBR are not always positively correlated. For example, for TPIA, the NBR for C/C++ is lower than that for Java, but C/C++ has a higher STF score. This indicates that there are more significant and effective attack examples in C/C++. While attacks on Java may be more comprehensive, they are harder to achieve stable and significant results. Besides, whether it is NBR or STF, TPIA significantly outperforms the baselines.
Stealthiness. We hereby evaluate the trade-off between effectiveness and stealthiness. As we discussed in Section III-C, our TPIA can circumvent existing malicious code detection methods because they only use non-functional code to achieve malicious goals. Arguably, human inspection is currently the only feasible approach to detect TPIA attacks. Therefore, we measure the stealthiness of TPIA by the number of tokens and chars (NIT and NIC), which directly determine whether the victim will be aware of the perturbation.
As shown in Figure 6, ASR and STF show a clear increase as NIT increases. Before reaching 60% ASR, TPIA attacks require very low NIT/NIC values. If an attacker aims for an effective but unstable attack, the perturbation can be extremely stealthy. When ASR exceeds 80%, the increase in NIT/NIC values accelerates, but increasing NIT leads to a very stable and effective improvement in ASR and STF (only when ASR is close to 100% will there be a small fluctuation). This demonstrates that enhancing TPIA’s performance does not require significantly increasing the length of the perturbation as with existing methods [22], adding just a few perturbation tokens can be highly effective.
Additionally, we find that after ASR reaches above 80%, the growth of NIC is slower compared to NIT, while this difference is not significant below 80% ASR. This is likely because when NIT is small, longer character-length tokens can convey more information, whereas when NIT is large, shorter character-length tokens can convey more precise information. The tokens used by TPIA do not have distinct characteristics in terms of character length but are determined by the attacker’s design. This means that filtering TPIA attacks based on character length is unreliable.
@@
Figure 6: The study of stealthiness. We plot the above two graphs using NIT and ASR as the x-axis respectively. We sampled 11 different NITs, NIT 0 (representing No Attack) and NIT 2 to 20 (with a sampling interval of 2).
V-C Evaluation on Commercial Code LLMs
In previous experiments, we have demonstrated that TPIA reaches sufficient effectiveness, flexibility, and stealthiness on open-source models. However, some widely used models are commercial and therefore are closed-source in real-world scenarios. In these cases, attackers need to conduct transfer attacks, i.e., generating perturbations using open-source Code LLMs to attack closed-source ones. We hereby exploit two representative real-world scenarios to evaluate TPIA’s transferability for further verifying its realistic threat.
@@ V-C Evaluation on Commercial Code LLMs
    Gray-Box Setting. The attacker can infer the open-source model used by the target application through interaction but cannot obtain the exact version of the model. For example, the scale of the model’s parameters, fine-tuning information, and version numbers are unknown to the attacker. We simulate this scenario using CodeGeeX2-6b and the CodeGeeX application. To the best of our knowledge, the CodeGeeX model has at least four different versions, and the CodeGeeX application does not disclose the specific model information it uses. This is a typical gray-box scenario.
@@
    Black-Box Setting. A more challenging situation is when the attacker knows nothing about the model used by the target application, or the model itself is closed-source (such as CodeX). We simulate this scenario by attacking GitHub Copilot (supported by CodeX). To align with gray-box setting, we adopt CodeGeeX2-6b to generate malicious perturbations.
TABLE IV: The evaluation of transferability. We test TPIA’s ASR, NBR, and STF under gray-box and black-box scenarios respectively to demonstrate its transferability.
@@
Figure 7: The example of a successful attack in a real-world black-box scenario (Github Copilot) across files. The perturbation is inserted into the comment of a function. When the user calls a function in a code context, Copilot will be induced to generate malicious code at the specified position (after TRG).
@@
Settings. We construct all cases in three languages within the integrated application of VSCode and manually test randomly selected code completion samples from the Humaneval/Humaneval-x dataset (620 data items in total). Other settings are the same as those described in Section V-B.
@@
Results. As shown in Table IV, in the gray-box scenario, TPIA still demonstrates great attack performance. Specifically, TPIA achieves ASRs of 80.4% and 93.3% in Python and C/C++, respectively. Even in the worst-performing case (i.e., Java), the ASR achieves 71.7%, with only a 9.1% decrease compared to the white-box scenario; In the black-box scenario, TPIA also achieves a satisfying attack performance of 90.4% on Python. The ASRs on C/C++ and Java are 67.7% and 74.4%, respectively. In other words, our TPIA still maintains a high level of threat, although its performance is lower than that of white-box scenarios to some extent. In addition, TPIA achieves a high NBR ($>0.7$) and STF ($>0.6$) in both the gray-box and black-box scenarios, indicating that it also maintains sufficient flexibility. Overall, TPIA achieves good attack performance in both gray-box and black-box scenarios, demonstrating its reliable transferability and threat to the commercial applications in the real-world.
@@
In Figure 7, we present a representative case of ST2 under the black-box scenario. The concise perturbation “TR os now rm sys” is inserted in the comment of a tool function. The commercial applications generate the functional code “os.system("rm -rf /")” at the attacker-specified position (i.e., after using TRG function) under the induction of the perturbation, although it is obviously a highly malicious instruction that can cause dangerous results.
V-D Ablation Study
In this section, we conduct an ablation study to demonstrate the necessity of the keyword-based designing and forward reasoning enhancement methods proposed in Section 4 for TPIA attacks. We also discuss the effects of a critical method-related hyper-parameter (i.e., Top-$k$), which is crucial for determining both the computational load of Algorithm 1 and the effectiveness of our TPIA.
(a) ASR Study
(b) STF Study
Figure 8: Ablation study of keywords number and Top-$k$. We compare TPIA’s ASR and STF under three conditions: without keywords, with 1-keyword, and with 2-keywords. During the process of generating perturbations, we use six different Top-$k$ values in our evaluation.
Effects of Top-$k$. This experiment validates the impact of increasing the Top value on the performance of TPIA. As shown in Figure 8, increasing the Top-$k$ value can significantly improve TPIA’s ASR and STF. Note that, the attack performance can reach its optimal level when the Top-$k$ is increased to the value of the vocab size of LLM. However, this approach would introduce computationally expensive overhead (CodeGemma has a vocab size of 256,000, far exceeding our default Top-$k$ value of 400). As such, the attacker should assign Top-$k$ value based on their specific needs.
(a) No FRE
(b) FRE
Figure 9: Ablation study of FRE. We compare the differences in ASR and loss function values both with and without using FRE module. 1-key1 and 1-key2 represent keeping one of the two selected keywords, respectively.
Effects of Keyword-based Designing. This experiment evaluates the effects of our optimization module, i.e., keyword-based designing. Specifically, we vary the number of keywords and generate the perturbations. As shown in Figure 8, limiting the number of keywords has a significant impact. The result of 1-keywords perturbation has approximately 10% decreases of ASRs and above 0.3 decrease of STF compared to 2-keywords perturbation. The 0-keyword perturbation, even with high computational power (600 Top-$k$ value), struggles to achieve an ASR above 80%. In contrast, 2-keywords perturbation can easily achieve both high ASR (above 95%) and STF (above 0.95). The above results indicate that using keywords can effectively improve the performance of TPIA.
Effects of Forward Reasoning Enhancement. We hereby perform a representative case study to illustrate the effects of forward reasoning enhancement (FRE), which corrects the distortion of the loss function. Specifically, we use the ‘ST2’ threat case as an example for our discussions since it is representative and difficult for attackers to achieve. It requires generating a clearly malicious code snippet and necessitates producing target code in a completely unrelated context, making it hard for the Code LLM to generate the target code, easily leading to distortion. To better evaluate the effects of this module, we first mitigate the effects of keyword-based designing by using static instead of optimized keywords. We judge the degree of distortion based on the loss value. The results are illustrated in Figure 9, where we have three key findings. Firstly, under the same conditions, using FRE can lead to a higher ASR. Secondly, without FRE, the ASR is still very low ($<20\%$) even when the loss value is lower than that with FRE. Thirdly, FRE has a greater impact on ASR than keyword-based design. These results verify the necessity of forward reasoning enhancement in our TPIA method.
VI Discussions
VI-A Resistance to Potential Defenses

### v6

@@ 4.1 Main Settings
TABLE II: The ASR across different Code LLMs. The experiment targets four code completion datasets and tests the success rates of various attack methods using three Code LLMs. As shown in Table I, some datasets do not include data for the C/C++ and Java languages, thus we have a total of 7 subdatasets. Best attack results are marked in boldface.
@@
Malicious Objectives and Code Datasets. We hereby collect 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) [27] across 3 programming languages: Python, Java, and C/C++. The straightforward objectives include using MODE_ECB for encryption (dubbed ‘ST0’), using ssl3 protocol (dubbed ‘ST1’), and executing “rm -rf /” (dubbed ‘ST2’). ST0 and ST1 are studied in classical code backdoor attacks [10], while ST2 is a famous destructive instruction. CWE vulnerabilities are generally selected for experiments in code model research [52, 15], and CWE Top-10 KEV weaknesses are supported by American cybersecurity and infrastructure security agency (CISA). We construct cases using the method described in Section 3.3 based on the official examples (i.e., CWE Top-10 KEV) or LLM-generated demo (i.e., ST0, ST1, ST2). As shown in Table I, some objectives apply only to C/C++, e.g., CWE416 “Use After Free”, thus we have a total of 31 cases. The details of these cases are discussed in Appendix A. In addition, we collect corresponding code completion datasets, including Humaneval [53], Humaneval-x [5], CodeXGLUE [54], MBPP [55], and Eval-Plus [56]. The items in these datasets contain prompts and corresponding outputs across different code completion tasks, which can simulate unpredictable users’ instructions and context in real-world scenarios. Some datasets do not cover Java or C/C++, as shown in Table I. As a result, we have 7 sub-datasets (i.e., the subsection of a dataset containing a particular programming language) containing more than 14,000 data items totally.
@@
Model Selection. In our experiment, we select three models from two representative Code LLM series, including CodeGemma [4] and CodeGeeX [5], for evaluation. CodeGemma series are trained by Google LLC and outperform some old-version Code LLMs like CodeLlama [31] on code completion tasks. We select CodeGemma-2b and CodeGemma-7b to compare the performance differences between models of distinct scales. CodeGeeX series have corresponding applications [2] with more than 935k downloads (on VScode [28] marketplace). We select CodeGeeX2-6b, a version with stable performance. We exclude earlier small-scale code-oriented models, such as CodeBERT [57] and CodeT5 [58], because large-scale code-oriented models perform much better than them in code completion tasks.
@@
Baseline Selection. We compare our ShadowCode to GCG-based and manually designed IPI attacks. In particular, the vanilla GCG [26] is a representative PI attack and does not meet the requirements of our attack scenarios. Therefore, we adapt GCG into an IPI attack that is applicable to Code LLMs, dubbed ‘C-GCG’ (code-oriented GCG). The manually designed IPI method (dubbed ‘IPI-M’) is designed based on the existing work [18]. IPI-M limits the number of injected tokens to make the perturbation concise since the existing IPI attack usually uses a large number of readable comments. Besides, we also provide the results of the original model (dubbed ‘No Attack’) for reference, since the original model itself may also have inherent insecurity factor. More details for baselines are provided in Appendix A.
@@
where $r_{1}$ and $r_{2}$ are set to 20% and 80% respectively in our evaluation empirically. Further reasons of evaluation metrics selection are discussed in Appendix A.
@@
Hyperparameters and Computational Facilities. We conduct experiments using the following parameter settings to evaluate our approach. During gradient calculation, we utilize five noisy code samples, and the random seed for initializing induced perturbations is set to 100. For forward reasoning enhancement module, the number of tokens reinforced is set to $h=2$, as defined in Equation (3). For the hyperparameters related to attack performance, the default parameter values are as follows: the Top-$k$ value is 400, the rate $r$ in Equation (3) is 0.4, the number of keyword is 2, and the perturbation length is 10. Tokenization to calculate token number is performed using Gemma’s tokenizer.
@@
We run Algorithm 1 on an NVIDIA RTX A6000 Graphics Card with the above parameter settings, employing a float16 precision format without quantization. In general, the perturbation generation process requires approximately 0.5-1 hours for CodeGemma-2b and 2-3 hours for both CodeGemma-7b and CodeGeeX-6b. These computational times ensure the generation of high-quality induced perturbations tailored to the respective model sizes across all the experiment threat cases.
4.2 Evaluation on Open-source Code LLMs
@@
We first perform the attack evaluation on open-source Code LLMs. This includes three aspects: (1) we assess the attack effectiveness over 3 Code LLMs across all threat cases, demonstrating that ShadowCode is widely effective and more threatening than other attacks; (2) we evaluate attack flexibility against different targets using two special metrics; (3) we evaluate the trade-off between effectiveness and stealthiness $w.r.t.$ the number of injected tokens.
@@
Performance on Effectiveness. We hereby evaluate whether ShadowCode can successfully induce the Code LLMs to generate attacker-specified targets. We control the number of injected tokens for three attack methods in this experiment. For ShadowCode, we use its default settings: perturbation length of 10 and 2 keyword, while C-GCG uses a prefix length of 12. Since IPI-M requires a sufficient number of tokens to be effective, we just limit its perturbation to fewer than 30 tokens. Table II provides the average ASR for four methods. The effectiveness of ShadowCode significantly outperforms the other baseline methods: it achieves an average ASR of over 80% across all three Code LLMs, with the highest average ASR reaching 86.6%.
@@
Comparing the three Code LLMs, the CodeGemma-7b model has the highest security in normal conditions (8.4% for No Attack) and is the least likely to generate vulnerable code. However, it is the most susceptible model to ShadowCode, which may be attributed to its strong contextual learning ability. Comparing the three programming languages, Python is the most susceptible language to ShadowCode due to its concise syntax and the smallest number of tokens consumed; Besides, ShadowCode is relatively less effective on CodeXGLUE, which is mostly due to its larger data length and more complex data content. In contrast, the simplest dataset with the shortest data items, MBPP, is the most vulnerable to attacks, including C-GCG and IPI-M. Therefore, we believe that an increase in the contextual complexity can reduce the ASR of ShadowCode. In other words, when the contextual code becomes complex, the strength of ShadowCode (e.g., the length of induced perturbations) needs to be increased accordingly to ensure the attack’s effectiveness. According to our subsequent experiments in Section 4.3 and Section 4.5, increasing the length of ShadowCode or the Top-$k$ value can effectively enhance the attack performance of ShadowCode. Additionally, we further enrich our experimental results by exploring the attack effectiveness for each case in Appendix B.
@@
TABLE III: The attack evaluation results of NBR and STF against open-source Code LLMs. Best results are marked in bold.
Metric Method C/C++ Python Java
  NBR No Attack 0.00 0.22 0.01
C-GCG 0.39 0.50 0.56
IPI-M 0.08 0.25 0.17
Ours 0.77 0.92 0.78
STF No Attack 0.00 0.11 0.03
C-GCG 0.26 0.47 0.39
IPI-M 0.09 0.19 0.07
Ours 0.76 0.87 0.74
@@
Flexibility. Previous experiments indicate that ShadowCode exhibits the best average ASR performance across various models. However, in real-world scenarios, attackers may have more complex demands. Generally, these demands can be categorized into two situations. One is when attackers want to achieve diverse malicious objectives, which requires that cases with low ASR be minimized. The other is when attackers want to cause severe damage, which requires that cases with high ASR be maximized. We refer to the capability to meet these complex demands as ”flexibility.” NBR is used to evaluate attack flexibility for the first situation, and STF for the second.
@@
As shown in Table III, ShadowCode has the highest NBR on the Python programming language, which means that implementing viable threat cases in Python is the easiest, and attacks targeting Python files are more flexible. Additionally, ShadowCode scores the highest in STF for Python, indicating that the attack effects in Python are more significant and pose a more serious real-world threat. STF and NBR are not always positively correlated. For example, for ShadowCode, the NBR for C/C++ is lower than that for Java, but C/C++ has a higher STF score. This indicates that there are more significant and effective attack examples in C/C++. While attacks on Java may be more comprehensive, they are harder to achieve stable and significant results. Besides, whether it is NBR or STF, ShadowCode significantly outperforms the baselines.
@@
Figure 6: The study of stealthiness. We plot the above two graphs using NIT and ASR as the x-axis respectively. We sampled 11 different NITs, NIT 0 (No Attack) and NIT 2 to 20 (with a sampling interval of 2).
Stealthiness. We hereby evaluate the trade-off between effectiveness and stealthiness. ShadowCode can circumvent existing malicious code detection methods because they only use non-functional code to achieve malicious goals. Arguably, human inspection is currently the only feasible approach to detect ShadowCode attacks. Therefore, we measure the stealthiness of ShadowCode by the number of tokens and chars (NIT and NIC), which directly determine whether the victim will be aware of the perturbation.
As shown in Figure 6, ASR and STF show a clear increase as NIT increases. Before reaching 60% ASR, ShadowCode attacks require very low NIT/NIC values. If an attacker aims for an effective but unstable attack, the perturbation can be extremely stealthy. When ASR exceeds 80%, the increase in NIT/NIC values accelerates, but increasing NIT leads to a very stable and effective improvement in ASR and STF (only when ASR is close to 100% will there be a small fluctuation). This demonstrates that enhancing ShadowCode’s performance does not require significantly increasing the length of the perturbation as with existing methods [18], every token can effectively improve the ASR.
Additionally, we find that after ASR reaches above 80%, the growth of NIC is slower compared to NIT, while this difference is not significant below 80% ASR. This is likely because when NIT is small, longer character-length tokens can convey more information, whereas when NIT is large, shorter character-length tokens can convey more precise information. The tokens used by ShadowCode do not have distinct characteristics in terms of character length but are determined by the attacker’s design of the token number. This means that filtering ShadowCode attacks based on character length is unreliable.
4.3 Evaluation on Commercial Code LLMs
In previous experiments, we demonstrate that ShadowCode reaches sufficient effectiveness, flexibility, and stealthiness on open-source models. However, some widely used models are commercial and therefore are closed-source in real-world scenarios. In these cases, attackers need to conduct transfer attacks, i.e., generating induced perturbations using open-source Code LLMs to attack closed-source ones. We hereby evaluate ShadowCode’s transferability by leveraging two representative real-world scenarios, thereby further verifying its practical threat.
@@ 4.3 Evaluation on Commercial Code LLMs
    Gray-Box Setting. The attacker can infer the open-source model used by the target application through interaction but cannot obtain the exact version of the model. For example, the scale of the model’s parameters, fine-tuning information, and version numbers are unknown to the attacker. We simulate this scenario using CodeGeeX2-6b and the CodeGeeX application. To the best of our knowledge, the CodeGeeX model has at least four different versions, and the CodeGeeX application does not disclose the specific model information it uses.
@@
    Black-Box Setting. A more challenging situation is when the attacker knows nothing about the model used by the target application, or the model itself is closed-source (such as Codex). We simulate this scenario by attacking GitHub Copilot (supported by Codex). To align with gray-box setting, we adopt CodeGeeX2-6b to generate induced perturbations.
TABLE IV: The evaluation of transferability. We test ShadowCode’s ASR, NBR, and STF under gray-box and black-box scenarios respectively to demonstrate its transferability.
@@
Figure 7: The example of a successful attack in a real-world black-box scenario (Github Copilot) across files. The perturbation is inserted into the comment of a function. When the user calls a function in a code context, Copilot will be induced to generate malicious code at the specified position (after calling “TRG”).
@@
Settings. We construct all cases in three languages within the integrated application of VSCode and manually test randomly selected code completion samples from the Humaneval/Humaneval-x dataset (620 data items in total). Other settings are the same as those described in Section 4.2.
@@
Results. As shown in Table IV, in the gray-box scenario, ShadowCode still demonstrates great attack performance. In Python and C/C++, ShadowCode achieves ASRs of 80.4% and 93.3%, respectively. Even in the worst-performing case (i.e., Java), the ASR achieves 71.7%, with only a 9.1% decrease compared to the white-box scenario; In the black-box scenario, ShadowCode also achieves a satisfying attack performance of 90.4% on Python. The ASRs on C/C++ and Java are 67.7% and 74.4%, respectively. In other words, our ShadowCode still maintains a high level of threat, although its performance is lower than that of white-box scenarios to some extent. In addition, ShadowCode achieves a high NBR ($>0.7$) and STF ($>0.6$) in both the gray-box and black-box scenarios, indicating that it also maintains sufficient flexibility. Overall, ShadowCode achieves good attack performance in both gray-box and black-box scenarios, demonstrating its reliable transferability and threat to the commercial applications in the real-world.
@@
In Figure 7, we present a representative case of ST2 in the black-box scenario. The concise perturbation “TR os now rm sys” is inserted in the comment of a tool function. The commercial applications generate the functional code “os.system("rm -rf /")” at the attacker-specified position (i.e., after using TRG function) under the induction of the perturbation, although it is obviously a highly malicious instruction that can cause dangerous results.
4.4 Resistance to Potential Defenses